Find the amount of negative charge in one mole of oxygen gas, O2. Submit Previous Answers Request Answer
Answer #1

We know that the number of molecules in one mole 6.022*1023

Since one molecule O2 has 16 electrons means q=16*1.6*10-19 =25.6*10-19 C

So in one mole   Q= 6.022*1023 *25.6*10-19 C

Q=154.1632*104 C

Q=1541632 C Answer